OpenStudy (anonymous):

Integration! \[\int \frac{x-3}{x^2 + 2x + 2} dx\]

OpenStudy (anonymous):

I couldn't see a u that works as-is. I'm tried completing the square to see if that did anything productive. \[\int \frac{x-3}{(x+1)^2 +1}dx\]\[u = x + 1 \\ du = dx\] So. \[\int \frac{u - 4}{u^2 + 1} du\] That's a little better.

zepdrix (zepdrix):

Hmm have you learned trig sub's yet? It works quite nicely for this problem :o

OpenStudy (anonymous):

I know them in theory. In practice I'm a bit shaky. Hold my hand and I'll give it a shot? :D

zepdrix (zepdrix):

Actually your u sub works our really well also,\[\Large \int\limits\limits \frac{u - 4}{u^2 + 1} du\quad=\quad \int\limits\limits \frac{u}{u^2+1}du-4\int\limits \color{orangered}{\frac{1}{u^2+1}}du\]

zepdrix (zepdrix):

Do you remember that `orange` integral by chance? :) It's a good one to memorize!

OpenStudy (anonymous):

No, but I have a table of integrals close at hand and that's one of those pretty ones that is surely on it. :D

zepdrix (zepdrix):

So we split it up into a couple of parts. For the first part, you can apply another substitution if it's not jumping out at you yet :) m=u^2+1 should get the job done.

OpenStudy (anonymous):

\[v = u^2 + 1 \\ dv = 2u du\]\[\frac{1}{2} \int \frac{1}{v} dv = \ln |v| = \ln |u^2 + 1| = \ln |(x+1)^2 + 1| + C\] Yes, that's not so bad. So combining that with ye olde integral table I get: \[\frac{\ln |x^2 +2x + 2|}{2} - 4\tan^{-1}(x + 1) + C\]

zepdrix (zepdrix):

Yay good job \c:/

OpenStudy (anonymous):

Out of curiosity, what would be the trig sub?

OpenStudy (anonymous):

I guess I can kind of work backwards from my answer. Something involving tan!

zepdrix (zepdrix):

After completing the square as you did,\[\Large \int\limits \frac{x-3}{(x+1)^2 +1}dx\]We could utilize a trig sub to get rid of the nasty addition in the denominator. We would let,\[\Large x+1\quad=\quad \tan \theta\]\[\Large x-3\quad=\quad \tan \theta-4\]\[\Large dx\quad=\quad \sec^2\theta\;d \theta\]Those are the pieces we would use for our substitution.

zepdrix (zepdrix):

\[\Large \int\limits\limits \frac{x-3}{(x+1)^2 +1}dx\qquad\to\qquad\int\limits\limits \frac{\tan \theta-4}{(\tan \theta)^2 +1}(\sec^2\theta \;d \theta)\]

zepdrix (zepdrix):

See how our trig identity will get rid of that nasty addition in the bottom? :)

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

\[\int \frac{\tan \theta - 4}{(\tan \theta)^2 + 1} \sec^2\theta d\theta\]\[\int \frac{\tan \theta - 4}{\sec^2 \theta} \sec^2\theta d\theta\]\[\int (\tan \theta -4) d\theta\]

zepdrix (zepdrix):

Yah! It simplifies down quite nicely huh? Tangent is another one of those weird integrals that you might want to just memorize. You can change it to \(\Large \dfrac{\sin\theta}{\cos\theta}\) and do a u-sub from there if you want though.

OpenStudy (anonymous):

Ok. So.\[v = \cos \theta \\ dv = -\sin \theta\] \[- \int \frac{1}{v} dv = -\ln |v| + C = -\ln |\cos \theta| + C\] Plus the \(4 \int d\theta\) leaves \(4\theta - \ln |\cos \theta| + C\) Ok, so now how to get x back. \[x + 1 = \tan \theta\]\[x + 1 = \frac{\sin \theta}{\cos \theta}\] Er.. I can solve that for \(\cos \theta\) but there is still a theta floating around.

OpenStudy (anonymous):

Oh, the 4 was minus. Stupid details. :D But anyway.

OpenStudy (anonymous):

I can see from comparing my two answers that somehow \[\frac{\ln |x^2 + 2x + 2|}{2} + C= -\ln |\cos(\tan^{-1}(x+1)))| + C\] But hopefully I can be forgiven for not grasping the immediate relationship.

zepdrix (zepdrix):

Oh yes there is another annoying step when doing trig subs. Undoing the substitution is a bit tricky. We have to go back to good ole triangle math! :)\[\Large \tan \theta\quad=\quad x+1\qquad=\quad\frac{x+1}{1}\quad=\quad\frac{opposite}{adjacent}\]

zepdrix (zepdrix):

We draw a triangle, label one of the angles theta, and then label the sides according to our relationship.

OpenStudy (anonymous):

Oh! How geometric.  So \(\cos \theta = \frac{1}{\sqrt{(x+1)^2 + 1}}\) Among other things. I think Pythagoras is making it worse?

zepdrix (zepdrix):

Worse or better .. I dunno. But it's leading us to the same answer we got using the other method! :O

OpenStudy (anonymous):

Oh! \[-\ln |(x^2 + 2x + 2)^{-1/2}|\]

zepdrix (zepdrix):

Yah there ya go! :D

OpenStudy (anonymous):

Now I see it! Thank you. :D

zepdrix (zepdrix):

yay team c:

OpenStudy (anonymous):

It's quite clear now. Thank you for the detailed explanations. I would give you another medal if I could. :P :D